OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 1 3x + 2y = 2

OpenStudy (anonymous):

need help

OpenStudy (sriram):

multiply first eq. by 2 and subtract we get x then substitute and solve y

OpenStudy (sriram):

answer would be (0,1)

OpenStudy (anonymous):

y = 1-5x 3x+2(1-5x)=2 3x+2-10x=2 -7x=0 x = 0 y = 1 i think